### v1

@@ Abstract
Large Language Models (LLMs) with long context capabilities are integral to complex tasks in natural language processing and computational biology, such as text generation and protein sequence analysis. However, training LLMs directly on extremely long contexts demands considerable GPU resources and increased memory, leading to higher costs and greater complexity. Alternative approaches that introduce long context capabilities via downstream finetuning or adaptations impose significant design limitations. In this paper, we propose Fully Pipelined Distributed Transformer (FPDT) for efficiently training long-context LLMs with extreme hardware efficiency. For GPT and Llama models, we achieve a 16x increase in sequence length that can be trained on the same hardware compared to current state-of-the-art solutions. With our dedicated sequence chunk pipeline design, we can now train 8B LLM with 2 million sequence length on only 4 GPUs, while also maintaining over $55\%$ of MFU. Our proposed FPDT is agnostic to existing training techniques and is proven to work efficiently across different LLM models. The code is available here.
@@ 1 Introduction
However, LLM training is typically constrained to relatively short context lengths, such as 8K or 32K tokens. Currently, most LLMs utilize rotary position embedding Su et al. (2024) (RoPE) to encode input tokens. Despite its improved efficiency and extrapolation capability compared to regular position embeddings Vaswani (2017), to accommodate much longer prompts during inference, RoPE often requires aggressive rescale and map. These adjustments struggle in properly adapt models to longer context before the model’s performance deteriorates, leading to a collapse in the quality of its outputs. This presents a critical challenge: the necessity to train LLMs originally on the desired long context lengths to ensure robust performance across varying applications.
@@ 2.1 Memory-efficient Transformer
The substantial memory demands of Transformers have spurred extensive research into memory-efficient attention mechanisms to facilitate their application to longer sequences. FlashAttention Dao et al. (2022); Dao (2023) has emerged as the de facto standard due to its robust and versatile performance. FlashAttention employs an online softmax technique to avoid the materialization of the attention matrix, thereby reducing memory overhead from $O(N^{2})$ to $O(N)$ while preserving accuracy. Other notable strategies include low-rank approximations Wang et al. (2020); Katharopoulos et al. (2020), kernel-based methods Kitaev et al. (2020); Xiong et al. (2021); Lu et al. (2021), and sparse attention mechanisms Child et al. (2019), which approximate or selectively compute attention to minimize memory consumption. Furthermore, techniques that combine local and global contexts Ainslie et al. (2020); Liu et al. (2021); Beltagy et al. (2020); Zaheer et al. (2020) enable superior performance on tasks involving long sequences or large-scale inputs while maintaining computational efficiency. Our work is inspired by FlashAttention and builds upon it by introducing hierarchical blockwise computation, leveraging the memory hierarchy in modern systems, which significantly enhances scalability and reduces memory requirements.
@@ 2.2 Long context training
Megatron-SPKorthikanti et al. (2023) adopts a sequence parallelism technique which is tightly integrated with its tensor parallelism. In this approach, sequences are partitioned along the sequence dimension, and all-gather and reduce-scatter collectives are employed to aggregate the QKV (query, key, value) projections for attention computation. The communication complexity analysis indicates that, in contrast to our approach, the communication volume in Megatron-SP’s sequence parallelism increases linearly with the message size (M) regardless of the number of compute devices.
@@
Ring Attention Liu et al. (2023) enhances Transformer’s scalability by distributing long sequences across multiple devices. This innovative approach overlaps the communication of key-value pairs with the computation of blockwise attention, effectively increasing the feasible sequence length proportionally to the number of available devices. However, reliance on device count for scaling and multi-step communications introduces potential issues in environments with sub-optimal hardware, where performance can be unpredictably affected by network latency and bandwidth constraints.
DeepSpeed Ulysses Jacobs et al. (2023) tackles the challenges of sequence parallelism by partitioning input data along the sequence dimension and utilizing an efficient all-to-all collective communication strategy for attention computations. Although this method maintains a constant communication volume regardless of the increase in sequence lengths and device counts, achieving significant speedups and scalability, it may still encounter practical hurdles in deployment related to large-scale clusters and the optimization of collective communication patterns across diverse computing environments.
Each of these methods uniquely contributes to the field of long-sequence Transformer training, offering solutions focused on memory efficiency, communication overhead, and computational speed. However, the common requirement for substantial GPU resources in all these approaches underscores a critical barrier to broader adoption and scalability. This challenge highlights the need for future research to develop more resource-efficient methods that can deliver similar benefits with a modest hardware budget.
@@ 3.1 GPU memory requirements in distributed Transformer
As Transformer becomes the de facto most powerful and ubiquitous architecture in today’s generative models, its overall pipeline also becomes homogeneous. We use the following table to show the required GPU memory at each step in forward and backward of a Transformer block.
Noticeable in this table, is that to get query, key, and value, the memory footprint is directly increased by three times, which solely can potentially lead to an out-of-memory issue when the sequence itself is too long to fit in the GPU memory. Also, as most Alltoall implementations do not support in-place operation, we need to create a receive buffer in addition to the input buffer. And if asynchronous communication is enabled, we need to prepare the receive buffer for query, key, and value at one time, leading to an $6Nd$ memory footprint.
@@ 4 Design of Fully Pipelined Distributed Transformer
In this section, we start by introducing how to design a seamless and efficient pipelined distributed attention with offloading. Since different hardware hierarchies are involved in the design, e.g. Tensor cores, NVLINK, PCIe, etc, each with different data throughput and latency, they are required to coordinate carefully.
@@ 4.1 Pipelining and scheduling
For the first QKV projection, since tokens are processed elementwise, we directly slice the local sequence tensor $[b,s_{local},h_{global},d]$ into $u$ chunks, each as a $[b,\frac{s_{local}}{u},h_{global},d]$ tensor. We denote these chunks as $T_{i}$, where $i\in\{0,1,\ldots,u-1\}$. As shown in Figure 4, $c_{i}$ is projected to query $q_{i}$, key $k_{i}$, and value $v_{i}$. Then, we perform the Alltoall communication among the sequence parallel group. Recall that in table 2, we mentioned that since Alltoall communication cannot be performed in-place, additional receive buffers need to be pre-allocated. In our chunk design, as we largely reduced the sequence length by a factor of $u$, even though the receiving buffers for $\hat{q_{i}},\hat{k_{i}},\hat{v_{i}}$ are still required, they become trivial compared to the original sequence tensor with full context.
@@
Figure 7: Double buffer leverages multiple CUDA streams in the backward pass of a Transformer block. We overlap most offloading operations with the attention gradients computation.
@@ 4.2 Double buffering
Figure 10: Average Time Spent in Alltoall, Attention Forward, Attention Backward, and three different host-to-device fetching strategies.
Thus, we care more about when the latency of attention computation overpasses that of host-to-device fetching. Unlike inter-GPU connections, data transfer among host memory and GPUs is not homogeneous. In our case, all GPUs will share the PCIe bandwidth, especially when host-to-device memory copy is issued simultaneously on each GPU. Therefore, we investigated two different strategies for fetching data from host memory to the corresponding GPUs. The most intuitive way is to let each GPU issue its own $HtoD$ transferring, which can make use of all DMA engines available, however, this might lead to resource contention as each GPU will try to consume some PCIe bandwidth as well as PCIe lanes. The second strategy is to let one GPU in each node fetch all related tensors, and then use a scatter operation to send corresponding chunks to other devices. In our profiling, we found that though the multi-GPU $HtoD$ strategy performs worse at smaller data sizes, due to the overhead in lane contention, latencies of both methods are overpassed by attention computation at around 32k to 64k, and thus, their difference becomes negligible as the sequence length grows larger. Since fetching from one GPU and scattering to other GPUs require additional synchronization and barrier, we choose the first strategy where we allow every GPU to fetch its data without any additional synchronization.
@@ 5.1 Experimental Setup
Experimental Environment: We use multiple GPU nodes, each with four A100 80 GB, connected via 3rd-Gen NVLink. There are two CPU sockets. The PCIe between host and device has a theoretical uni-directional bandwidth of 32 GB/s. Each node is equipped with 1 TB of host memory. For the internode connection, we use NVIDIA 200 Gbps HDR InfiniBand.
@@ 5.2 Overall Performance
There are several widely used solutions for training long-context language models. Megatron-SP Korthikanti et al. (2023) partitions sequence activations and leverages tensor parallel. DeepSpeed Ulysses Jacobs et al. (2023) adopts a one-step Alltoall to gather tokens and scatter heads among all GPUs. In this section, We compare our proposed design with these state-of-the-art solutions.
We choose six widely used LLM models with different sizes. For GPT-like ones, we have 2.7B, 6.7B, 13B, and 30B. For Llama, we use the 8B and 70B models. For Megatron-SP, we follow its default parallel setting, which leverages tensor model parallel and sequence parallel. For our proposed FPDT, ZeRO-3 model parameter sharding is leveraged, this ensures fair comparison between Megatron-SP, DeepSpeed Ulysses, and FPDT. For all experiments, we set the batch size to 1 and enable the activation checkpoint with CPU offloading. Our goal is to fully exploit the long sequence handling capability of each parallel strategy.
@@
Table 11 shows the end-to-end training performance. When running within one compute node, Megatron-SP and Ulysses exhibit similar hardware efficiency. For example, on the 2.7B GPT model, both methods have a maximum sequence length of 256K. For our proposed FPDT, with only chunking, we increase the sequence length by 8x longer, from 256K to 2M, without sacrificing performance. For the 6.7B GPT model, Megatron-SP and Ulysses support 128K and 256K sequence lengths. Our FPDT with chunking increases the bound to 512K while facing out-of-memory (OOM) issues when we keep increasing the sequence length. We then leverage the host memory. By offloading idle tokens from HBM, we support the 2M sequence length, which is 8x longer than the existing solutions. By properly managing the offloading, fetching, and computation, our FPDT with double buffer strategy reaches comparable hardware MFU as the non-offloading counterparts (i.e. FPDT w. chunking). Similar results are also shown in the 8B Llama 3 model training. As we increase the model’s size, we scale it up to multiple nodes. We found that Ulysses is generally more efficient than Megatron-SP, as the latter’s performance degrades severely when inter-node communication is included. For the 13B, 30B, and 70B models, we use 2, 4, and 8 nodes, respectively. While both Megatron-SP and Ulysses struggle with the memory spikes of activations, our FPDT works effectively by increasing the maximum sequence length to 2M while maintaining high MFU.
@@ 5.3 Tradeoff on Sequence Chunk Size
As discussed in  4.2, choosing a proper chunk size can not only exploit the computing power of the hardware but also allow the data moving from host to device and from device to host to overlap by computation. In Table 11, we use a default chunk size of 64K for all our FPDT-based methods. In this part, we will demonstrate why this is our preference. Table 12 shows the usage of GPU HBM at different chunk lengths. We conduct the profiling on a single node with 4 GPUs, with a fixed global sequence length of 256K. We change the length of each sequence chunk and compare the corresponding memory footprint and MFU. The chunk size of 256K means that we do not use chunk, but just run the baseline Ulysses. 8K, 16K, 32K, 64K, and 128K of chunk size corresponds to 32, 16, 8, 4, and 2 chunks in total in our FPDT pipeline scheme. As shown in Figure 12(a),  12(b),  12(d), our FPDT can significantly reduce the activation memory footprint; for example, in the 2.7B model, we reduce the activation memory from 27GB to 18GB by splitting the sequence into two chunks. Despite that increasing the number of chunks can further reduce the memory footprint, we found that 64K is a sweet point where the latency of offloading and prefetching can be hidden by the computation. Less chunks means a shorter pipeline, which makes latencies of the first data preparation and the last computation more salient, as they cannot overlap with other operations. More chunks, however, make the latency of computation too short to hide that of data loading. Therefore, we choose 64K as it has an overall best-overlapped pipeline.
@@ 5.5 Training strategies in long-context LLM
Though we discussed our experiments’ setup in 5.2, we would like to know how each strategy contributes to the long-context LLM training. Tensor parallel is widely used in distributed model training. It allows each GPU to only keep a slice of the tensor along the hidden dimension, hence also parallelizing the computation. However, it cannot reduce the memory footprint of intermediate buffers, since the GEMM between tensor [$N,B,C$] and weight [$\hat{C},C$] generates an intermediate buffer [$N,B,\hat{C}$] regardless of $C$. Activation checkpoint (AC.) is also a commonly used strategy in large model training, as it can significantly reduce the GPU memory pressure for models with many layers. Activation checkpoint with offloading (OC.) is a technique introduced in DeepSpeed, which allows the checkpointed activation to be moved to host memory. Both AC. and OC. play critical roles in enlarging the sequence length, i.e., from 32K to 128K to 512K. For Ulysses, it can seamlessly work with the ZeRO family, which reduces a decent amount of memory footprint, i.e., from 58.9G to 52.3G. However, ZeRO doesn’t mitigate the memory requirement for long-context LLM training as it reduces only memory usage related to the model parameters. Finally, our proposed FPDT, together with activation checkpointing, activation offloading, and ZeRO-3, remarkably increases the maximum sequence length by 8x, from 512K to 4M, on an 8B Llama 3 with 8 GPUs, while achieving extreme hardware efficiency.
5.6 Convergence evaluation
Figure 14: Loss curve in pretraining GPT models with 4 GPUs.
Figure 14 shows the convergence of the baseline GPT model that leverages tensor parallel on 4 GPUs, with a batch size of 256 and ZeRO-1 enabled, and our FPDT w/ and w/o offloading. Our proposed FPDT is a pure system optimization technique that enables the training of ultra-long sequence Transformer model, thus there is no (negative) on the quality of trained models.

### v2

@@ Abstract
Large Language Models (LLMs) with long context capabilities are integral to complex tasks in natural language processing and computational biology, such as text generation and protein sequence analysis. However, training LLMs directly on extremely long contexts demands considerable GPU resources and increased memory, leading to higher costs and greater complexity. Alternative approaches that introduce long context capabilities via downstream finetuning or adaptations impose significant design limitations. In this paper, we propose Fully Pipelined Distributed Transformer (FPDT) for efficiently training long-context LLMs with outstanding hardware efficiency. For GPT and Llama models, we achieve a 16x increase in sequence length that can be trained on the same hardware compared to current state-of-the-art solutions. With our dedicated sequence chunk pipeline design, we can now train 8B LLM with 2 million sequence length on only 4 GPUs, while also maintaining over $55\%$ of MFU. Our proposed FPDT is agnostic to existing training techniques and is proven to work efficiently across different LLM models. Code is available.
@@
Preliminary work. Under review by the Machine Learning and Systems (MLSys) Conference. Do not distribute.
@@ 1 Introduction
However, LLM training is typically constrained to relatively short context lengths, such as 8K or 32K tokens. Currently, most LLMs utilize rotary position embedding Su et al. (2024) (RoPE) to encode input tokens. Despite its improved efficiency and extrapolation capability compared to regular position embeddings Vaswani (2017), to accommodate much longer prompts during inference, RoPE often requires aggressive rescale and map. These adjustments struggle to properly adapt models to a longer context before the model’s performance deteriorates, leading to a collapse in the quality of its outputs. This presents a critical challenge: the necessity to train LLMs originally on the desired long context lengths to ensure robust performance across varying applications.
@@ 2.1 Memory-efficient Transformer
The substantial memory demands of Transformers have spurred extensive research into memory-efficient attention mechanisms to facilitate their application to longer sequences. FlashAttention Dao et al. (2022) utilizes an online softmax computation technique Milakov & Gimelshein (2018) and reduces memory requirements of self-attention from $O(n^{2})$ to $O(n)$ Rabe & Staats (2021) while preserving the accuracy. Other notable strategies include low-rank approximations Wang et al. (2020); Katharopoulos et al. (2020), kernel-based methods Kitaev et al. (2020); Xiong et al. (2021); Lu et al. (2021), and sparse attention mechanisms Child et al. (2019), which approximate or selectively compute attention to minimize memory consumption. Furthermore, techniques that combine local and global contexts Ainslie et al. (2020); Liu et al. (2021); Beltagy et al. (2020); Zaheer et al. (2020) enable superior performance on tasks involving long sequences or large-scale inputs while maintaining computational efficiency. Our work is inspired by FlashAttention and builds upon it by introducing hierarchical blockwise computation, leveraging the memory hierarchy in modern systems, which significantly enhances scalability and reduces memory requirements.
@@ 2.2 Long context training
Megatron-SPKorthikanti et al. (2023) adopts a sequence parallelism technique which is tightly integrated with its tensor parallelism. In this approach, sequences are partitioned along the sequence dimension, and all-gather and reduce-scatter collectives are employed to aggregate the QKV (query, key, value) projections for attention computation. The communication complexity analysis indicates that, in contrast to our approach, the communication volume in Megatron-SP’s sequence parallelism increases linearly with the sequence length regardless of the number of compute devices.
@@
Ring Attention Liu et al. (2023) enhances Transformer’s scalability by distributing long sequences across multiple devices. This innovative approach overlaps the communication of key-value pairs with the computation of blockwise attention, effectively increasing the feasible sequence length proportionally to the number of available devices. However, reliance on device count for scaling and multi-step communications introduces potential issues in environments with sub-optimal hardware regards network interconnects, where performance can be unpredictably affected by network latency and bandwidth constraints.
DeepSpeed Ulysses Jacobs et al. (2023) tackles the challenges of sequence parallelism by partitioning input data along the sequence dimension and utilizing an efficient all-to-all collective communication strategy for attention computations. Although this method maintains a constant communication volume regardless of the increase in sequence lengths and device counts, achieving significant speedups and scalability, it may still encounter practical hurdles in deployment related to large-scale clusters and the optimization of communication patterns across diverse computing environments.
Each of these methods uniquely contributes to the field of long-sequence Transformer training, offering solutions focused on memory efficiency, communication overhead, and computational speed. However, the common requirement for substantial GPU resources in all these approaches underscores a critical barrier to broader adoption in scientific research and deployment. This challenge highlights the need for future research to develop more resource-efficient methods that can deliver similar benefits with a modest hardware budget.
@@ 3.1 GPU memory requirements in distributed Transformer
As Transformer becomes the de facto most powerful and ubiquitous architecture in today’s generative models, its overall pipeline also becomes similar. Table 2 shows the required GPU memory at each step in the forward and backward of a Transformer block using DeepSpeed Ulysses.
Noticeable in this table, is that to get query, key, and value, the memory footprint is directly increased by three times, which solely can potentially lead to an out-of-memory issue when the sequence itself is too long to fit in the GPU memory. Also, as most Alltoall implementations do not support in-place operation, we need to create a receive buffer in addition to the input buffer. And if asynchronous communication is enabled, we need to prepare the receive buffer for query, key, and value at one time, leading to a $6Nd$ memory footprint, where $d$ denotes the hidden dimension.
@@ 4 Design of Fully Pipelined Distributed Transformer
In this section, we start by introducing how to design a seamless and efficient pipelined distributed attention leveraging both host and GPU memory. Since different hardware hierarchies are involved in the design, e.g. Tensor cores, NVLINK, PCIe, etc, each with different data throughput and latency, they are required to coordinate carefully.
@@ 4.1 Pipelining and scheduling
For the first QKV projection, since tokens are processed elementwise, we directly slice the local sequence tensor $[b,s_{local},h_{global},d]$ into $u$ chunks, each as a $[b,\frac{s_{local}}{u},h_{global},d]$ tensor. We denote these chunks as $T_{i}$, where $i\in\{0,1,\ldots,u-1\}$. As shown in Figure 4, $T_{i}$ is projected to query $q_{i}$, key $k_{i}$, and value $v_{i}$. Then, we perform the Alltoall communication among the sequence parallel group. Recall that in table 2, we mentioned that since Alltoall communication cannot be performed in-place, additional receive buffers need to be pre-allocated. In our chunk design, as we largely reduced the sequence length by a factor of $u$, even though the receiving buffers for $\hat{q_{i}},\hat{k_{i}},\hat{v_{i}}$ are still required, they become trivial compared to the original sequence tensor with the full context.
@@
Figure 7: Double buffer leverages multiple CUDA streams in the backward pass of a Transformer block. We overlap most offloading operations with the attention gradients computation. We only show attention-related operations for clarity.
@@ 4.2 Double buffering
Thus, we care more about when the latency of attention computation overpasses that of host-to-device fetching. Unlike inter-GPU connections, data transfer among host memory and GPUs is not homogeneous. In our case, all GPUs will share the PCIe bandwidth, especially when host-to-device memory copy is issued simultaneously on each GPU.
Figure 10: Average Time Spent in Alltoall, Attention Forward, Attention Backward, and two different host-to-device strategies.
Therefore, we investigated two different strategies for fetching data from host memory to the corresponding GPUs. The most intuitive way is to let each GPU issue its own $HtoD$ transferring, which can make use of all DMA engines available, however, this might lead to resource contention as each GPU will try to consume some PCIe bandwidth as well as PCIe lanes. The second strategy is to let one GPU in each node fetch all related tensors, and then use a scatter operation to send corresponding chunks to other devices. In our profiling, we found that though the multi-GPU $HtoD$ strategy performs worse at smaller data sizes, due to the overhead in lane contention, latencies of both methods are overpassed by attention computation at around 32k to 64k, and thus, their difference becomes negligible as the sequence length grows larger. Since fetching from one GPU and scattering to other GPUs require additional synchronization and barrier, we choose the first strategy where we allow every GPU to fetch its data without any additional synchronization.
@@ 5.1 Experimental Setup
Experimental Environment: We use multiple GPU nodes, each with four A100 80 GB, connected via 3rd-Gen NVLink. There are two CPU sockets. The PCIe between host and device has a theoretical unidirectional bandwidth of 32 GB/s. Each node is equipped with 1 TB of host memory. For the internode connection, we use NVIDIA 200 Gbps HDR InfiniBand.
@@ 5.2 Overall Performance
There are several widely used solutions for training long-context language models. Megatron-SP Korthikanti et al. (2023) partitions sequence activations and leverages tensor parallel. DeepSpeed Ulysses Jacobs et al. (2023) adopts a one-step Alltoall to gather tokens and scatter heads among all GPUs. Ring Attention leverages the online attention mechanism to distribute KV chunks and overlap the ring communication with local attention computation. However, the only end-to-end training framework with Ring Attention is Megatron-LM, which does not support activation checkpointing with CPU offloading. As we will also show in Table 3, this severely limits the maximum sequence length it supports. Thus, in this section, we compare our proposed design with the first two state-of-the-art solutions, both supported in DeepSpeed.
We choose six widely used LLM models with different sizes. For GPT-like models, we have 2.7B, 6.7B, 13B, and 30B. For Llama, we use the 8B and 70B models. For Megatron-SP, we follow its default parallel setting, which leverages tensor model parallel and sequence parallel. For our proposed FPDT, ZeRO-3 model parameter sharding is leveraged, this ensures fair comparison between Megatron-SP, DeepSpeed Ulysses, and FPDT. For all experiments, we set the batch size to 1 and enable the activation checkpoint with CPU offloading. Our goal is to fully exploit the long sequence handling capability of each parallel strategy.
@@
Figure 11 shows the end-to-end training performance. When running within one compute node, Megatron-SP and Ulysses exhibit similar hardware efficiency. For example, on the 2.7B GPT model, both methods have a maximum sequence length of 256K. For our proposed FPDT, with only chunking, we increase the sequence length by 8x longer, from 256K to 2M, without sacrificing performance. For the 6.7B GPT model, Megatron-SP and Ulysses support 128K and 256K sequence lengths. Our FPDT with chunking increases the bound to 512K while facing out-of-memory (OOM) issues when we keep increasing the sequence length. We then leverage the host memory. By offloading idle tokens from HBM, we support the 2M sequence length, which is 8x longer than the existing solutions. By properly managing the offloading, fetching, and computation, our FPDT with double buffer strategy reaches comparable hardware MFU as the non-offloading counterparts (i.e. FPDT w. chunking). Similar results are also shown in the 8B Llama 3 model training. As we increase the model’s size, we scale it up to multiple nodes. We found that Ulysses is generally more efficient than Megatron-SP, as the latter’s performance degrades severely when inter-node communication is included. For the 13B, 30B, and 70B models, we use 2, 4, and 8 nodes, respectively. While both Megatron-SP and Ulysses struggle with the memory spikes of activations, our FPDT works effectively by increasing the maximum sequence length to 2M while maintaining high MFU.
@@ 5.3 Tradeoff on Sequence Chunk Size
As discussed in  4.2, choosing a proper chunk size can not only exploit the computing power of the hardware but also allow the data moving from host to device and from device to host to overlap by computation. In Figure 11, we use a default chunk size of 64K for all our FPDT-based methods. In this part, we will demonstrate why this is our preference. Figure 12 shows the usage of GPU HBM at different chunk lengths. We conduct the profiling on a single node with 4 GPUs, with a fixed global sequence length of 256K. We change the length of each sequence chunk and compare the corresponding memory footprint and MFU. The chunk size of 256K means that we do not use chunk, but just run the baseline Ulysses. 8K, 16K, 32K, 64K, and 128K of chunk size corresponds to 32, 16, 8, 4, and 2 chunks in total in our FPDT pipeline scheme. As shown in Figure 12(a),  12(b),  12(d), our FPDT can significantly reduce the activation memory footprint; for example, in the 2.7B model, we reduce the activation memory from 27GB to 18GB by splitting the sequence into two chunks. Despite that increasing the number of chunks can further reduce the memory footprint, we found that 64K is a sweet point where the latency of offloading and prefetching can be hidden by the computation. Less chunks means a shorter pipeline, which makes latencies of the first data preparation and the last computation more salient, as they cannot overlap with other operations. More chunks, however, make the latency of computation too short to hide that of data loading. Therefore, we choose 64K as it has an overall best-overlapped pipeline.
@@ 5.5 Training strategies in long-context LLM
Though we discussed our experiments setup in 5.2, we would like to know how each strategy contributes to the long-context LLM training. Tensor parallel is widely used in distributed model training. It allows each GPU to only keep a slice of the tensor along the hidden dimension, hence also parallelizing the computation. However, it cannot reduce the memory footprint of intermediate buffers, since the GEMM between tensor [$N,B,C$] and weight [$\hat{C},C$] generates an intermediate buffer [$N,B,\hat{C}$] regardless of $C$. Activation checkpoint (AC.) is also a commonly used strategy in large model training, as it can significantly reduce the GPU memory pressure for models with many layers. Activation checkpoint with offloading (OC.) is a technique introduced in DeepSpeed, which allows the checkpointed activation to be moved to host memory. Both AC. and OC. play critical roles in enlarging the sequence length, i.e., from 32K to 128K to 512K. For Ulysses, it can seamlessly work with the ZeRO family, which reduces a decent amount of memory footprint, i.e., from 58.9G to 52.3G. However, ZeRO doesn’t mitigate the memory requirement for long-context LLM training as it reduces only memory usage related to the model parameters. Finally, our proposed FPDT, together with activation checkpointing, activation offloading, and ZeRO-3, remarkably increases the maximum sequence length by 8x, from 512K to 4M, on an 8B Llama 3 with 8 GPUs, while achieving extreme hardware efficiency.
5.6 Sparse attention
Sparse attention is crucial for improving the efficiency of long-context large language models (LLMs) by reducing the computational and memory costs of standard attention mechanisms, which scale quadratically with sequence length. By focusing only on the most relevant tokens, sparse attention significantly speeds up inference while preserving model accuracy. In Table 4, we use the block sparse attention to test the training MFU with our FPDT design. Note that, with sparse attention, only part of the tokens in key and value will be fetched from the host memory, while the query will always be the entire sequence. As we discussed in Figure 10, this may leads to an overall shift towards left, where the latency of fetching cannot be perfectly overlapped by computation, which explains the slightly lower MFU when sparsity is high.
Models Attention Sparsity
0.5 0.4 0.3 0.2 0.1 0.0
2.7B 41.7 38.6 38.8 39.0 38.9 38.4
8B 40.6 42.0 44.0 46.4 46.5 47.6
13B 40.7 42.8 44.2 45.4 46.4 46.1
Table 4: MFU(%) at different attention sparsity. A higher sparsity means less KV tokens are fetched for computation. We use 1 GPU for 2.7B model, and 4 GPUs for 8B and 13B models. Chunk size is 64K and global sequence length is 256K. 0.0 means full attention.